# Euler systems and relative Satake isomorphism

### v1

@@ 3.3.2. Homogeneous case
Let $\delta_{(X)}$ (resp. $\delta_{P(X)}$) be the modulus character for the good Levi subgroup $\mathbf{L}(\mathbf{X})$ (resp. parabolic subgroup $\mathbf{P}(X)$), cf. §2.1. The translate $\delta_{(X)}^{\frac{1}{2}}A_{X}^{*}$ defines a subvariety of $A^{*}$. Its ring of polynomial functions will be denoted by $\mathbb{C}[\delta_{(X)}^{\frac{1}{2}}A_{X}^{*}]$. We make the following canonical identification
(5) $\mathbb{C}[\Lambda_{X}]\xrightarrow{\,\sim\,}\mathbb{C}[A_{X}^{*}]\xrightarrow{\,\sim\,}\mathbb{C}[\delta_{(X)}^{\frac{1}{2}}A_{X}^{*}]$
@@
    Let $C_{c}^{\infty}(K\backslash G/K,\mathbb{C})\simeq\mathbb{C}[A^{*}]^{W}$ be the classical Satake isomorphism, then the above isomorphism is equivariant with respect to the natural action on the left hand side and multiplication on the right hand side.
Let $\mathcal{L}\in C_{c}^{\infty}(K\backslash G/K,\mathbb{C})$ be a Hecke operator, with Satake transform ^$\hat{\mathcal{L}}\in\mathbb{C}[A^{*}]^{W}$. By restriction and using identification (5), this defines an element
(6) ^$\hat{\mathcal{L}}_{(X)}\in\mathbb{C}[\delta_{(X)}^{\frac{1}{2}}A_{X}^{*}]\xrightarrow{\,\sim\,}\mathbb{C}[\Lambda_{X}]$
We would like to compute the function $\mathcal{L}\cdot\Phi_{0}$ in terms of $\hat{\mathcal{L}}$, which can be done using the inverse Satake isomorphism [Sak18]. To state it, recall that Sakellaridis defined a multiset $\Theta_{X}^{+}$ in [Sak13, §7.1]. By our split hypothesis, all the signs $\sigma_{\check{\theta}}=1$ in the notation of loc. cit., and the remaining two pieces of data can be combined as a representation of $\check{\mathbf{A}}\times\mathbb{G}_{m}$, cf. [BZSV24, §9.3.5].
@@
This is proven in [Sak18] under the additional assumption that $\mathbf{X}$ is wavefront. However, the only time the wavefront hypothesis is used in the derivation is to compute the normalization factor in the Plancherel measure. This exact normalization is not needed for the above result since the expressions do not depend on the choice of the $G$-invariant measure on $X$. ∎
@@ 3.3.3. Strongly tempered case
In this case, we have $\mathbf{A}=\mathbf{A}_{X}$, so we can drop the subscript $X$ from all combinatorial data. Moreover, $\delta_{(X)}=1$ and $\delta_{P(X)}=\delta$. Finally, all elements of $\Theta_{X}^{+}$ are in degree 1 by the recipe constructing them. Therefore, the following is the exact analogue of Theorem 3.6, with the final statement playing the role of [Sak13, Statement 7.1.5].
@@
3.4. Integral structure
Let the assumptions be as in the previous subsection. We give an ad hoc definition of an integral structure on $\mathbb{C}[\delta_{(X)}^{\frac{1}{2}}A_{X}^{*}]$, whose only purpose is to specify when a half power of $q$ can show up. This is abstractly explained by the analytic shearing construction in [BZSV24, §6.8]. Let $\mathbb{Z}_{(q)}=\mathbb{Z}[q^{-1}]$.
@@ 3.4. Integral structure
[…]
@@
where $a_{\check{\lambda}}\in\mathbb{Z}_{(q)}$. We are again identifying $\mathbb{C}[[A^{*}]]$ with $\mathbb{C}[[\Lambda]]$ by (1).
[…]
@@
[…]
@@
[…]
$\phi(x_{\check{\lambda}})$ $\in\delta_{P(X)}^{-\frac{1}{2}}(x_{\check{\lambda}})\delta_{(X)}^{\frac{1}{2}}(x_{\check{\lambda}})q^{\langle\check{\lambda},\rho\rangle}\mathbb{Z}_{(q)}$
$=q^{\langle\check{\lambda},-\rho_{P(X)}+\rho_{(X)}+\rho\rangle}\mathbb{Z}_{(q)}$
Since $\rho=\rho_{P(X)}+\rho_{(X)}$ and $2\rho_{P(X)}$ is integral, this is an integral power of $q$.
From this computation, we see that $\phi(x_{\check{\lambda}})$ is a polynomial in $\mathbb{Z}[q^{\pm 1}]$, where we now treat $q$ as a formal variable. To prove the required divisibility, we specialize it to $q=1$, so $\phi(x_{\check{\lambda}})|_{q=1}$ is the coefficient of $e^{\check{\lambda}}$ in the power series expansion of
^$\hat{\mathcal{L}}_{(X)}|_{q=1}\cdot\frac{\mathtt{ad}}{\mathtt{L}},\quad\text{where }\mathtt{ad}=\prod_{\check{\gamma}\in\check{\Phi}_{X}^{+}}(1-e^{\check{\gamma}}),\quad\mathtt{L}=\prod_{\check{\theta}\in\Theta^{+}}(1-e^{\check{\theta}}).$
Suppose $\gamma$ is a spherical root of type T, with corresponding coroot $\check{\gamma}$. Let $w_{\gamma}\in W_{X}$ be its associated simple reflection. We consider the behaviour of the above expression under $w_{\gamma}$. There are three terms.
  1. (1)
    The term ^$\hat{\mathcal{L}}_{(X)}|_{q=1}$ is $w_{\gamma}$-invariant by construction.
@@
    We need to understand the difference between $w_{\gamma}\Theta^{+}$ and $\Theta^{+}$. In the homogeneous case, this is described in [Sak13, Statement 7.1.5]. Since $\gamma$ has type T, there are two virtual colors $D$ and $D^{\prime}$ belonging to $\gamma$. Their associated valuations $v_{D}$ and $v_{D^{\prime}}$ sum to $\check{\gamma}$, so
    $\frac{w_{\gamma}\mathtt{L}}{\mathtt{L}}$ $=\frac{1-e^{-v_{D}}}{1-e^{v_{D}}}\frac{1-e^{-v_{D^{\prime}}}}{1-e^{v_{D^{\prime}}}}$
    $=e^{-(v_{D}+v_{D^{\prime}})}=e^{-\check{\gamma}}$
    In the strongly tempered case, the above discussion still holds due to our minuscule hypothesis combined with [SW22, Theorem 7.1.9(iii)].
@@
Performing the above computation for a spherical root of type G recovers equation (7), except with a plus sign. Therefore, we gain no information. In particular, all roots are of type G in the group case $\mathbf{H}\backslash\mathbf{H}\times\mathbf{H}$, so we do not see any additional divisibility in the classical Satake isomorphism.

### v2

@@ 3.3.2. Homogeneous case
Let $\delta_{(X)}$ (resp. $\delta_{P(X)}$) be the modulus character for $\mathbf{L}(\mathbf{X})\cap\mathbf{B}$ (resp. parabolic subgroup $\mathbf{P}(X)$), cf. §2.1. The translate $\delta_{(X)}^{\frac{1}{2}}A_{X}^{*}$ defines a subvariety of $A^{*}$. Its ring of polynomial functions will be denoted by $\mathbb{C}[\delta_{(X)}^{\frac{1}{2}}A_{X}^{*}]$. We make the following canonical identifications
(5) $\mathbb{C}[\Lambda_{X}]\xrightarrow{\,\sim\,}\mathbb{C}[A_{X}^{*}]\xrightarrow{\,\sim\,}\mathbb{C}[\delta_{(X)}^{\frac{1}{2}}A_{X}^{*}],$
@@
[…]
(6) $\hat{\mathcal{L}}_{(X)}\in\mathbb{C}[\delta_{(X)}^{\frac{1}{2}}A_{X}^{*}]\xrightarrow{\,\sim\,}\mathbb{C}[\Lambda_{X}]$
We would like to compute the function $\mathcal{L}\cdot\Phi_{0}$ in terms of $\hat{\mathcal{L}}$, which can be done using the inverse Satake isomorphism [Sak18]. To state it, recall that Sakellaridis defined a multiset $\Theta_{X}^{+}$ in [Sak13, §7.1]. By our split hypothesis, all the signs $\sigma_{\check{\theta}}=1$ in the notation of loc. cit., and the remaining two pieces of data can be combined as a representation of $\check{\mathbf{A}}\times\mathbb{G}_{m}$, cf. [BZSV24, §9.3.5]. We will write an element of $\Theta_{X}^{+}$ in the form $(\check{\theta},d_{\check{\theta}})$, where $\check{\theta}$ is a cocharacter, and $d_{\check{\theta}}\in\mathbb{Z}$.
@@
Remark 3.7.
We make an important comment regarding the interpretation of the right hand side. It is a rational function, and we need to expand it as a formal Laurent series before taking the restriction to $\Lambda_{X}^{+}$. To make sense of this expansion, one needs to specify a cone $\mathcal{C}_{X}\subseteq\Lambda_{X}$. In the above theorem, this is the strictly convex cone (positive-)dual to the highest weights appearing in the $\mathbf{G}$-module $F[\mathbf{X}]$, cf. [Sak18, §6]. In the homogeneous case, this is disjoint from $\Lambda_{X}^{+}$, so the right hand side describes a compactly supported function on $\Lambda_{X}^{+}$.
The cone will not be present in our notations since our proofs do not depend on its choice, but it is nonetheless important to note that there is a fixed choice.
@@
The result is [Sak18, Theorem 7.7], where it was proven under the additional assumption that $\mathbf{X}$ is wavefront. However, the only time the wavefront hypothesis is used in the derivation is to compute the normalization factor in the Plancherel measure. This exact normalization is not needed for the above result since the expressions do not depend on the choice of the $G$-invariant measure on $X$. ∎
@@ 3.3.3. Strongly tempered case
In this case, we have $\mathbf{A}=\mathbf{A}_{X}$, so we can drop the subscript $X$ from all combinatorial data. Moreover, $\delta_{(X)}=1$ and $\delta_{P(X)}=\delta$. Finally, all elements of $\Theta_{X}^{+}$ have $d_{\check{\theta}}=1$ by the recipe constructing them. Therefore, the following is the exact analogue of Theorem 3.6, with the final statement playing the role of [Sak13, Statement 7.1.5].
@@
3.3.4. Integral structure
Let the assumptions be as in the previous subsection. We give an ad hoc definition of an integral structure on $\mathbb{C}[\delta_{(X)}^{\frac{1}{2}}A_{X}^{*}]$, whose only purpose is to specify when a half power of $q$ can show up. This is abstractly explained by the analytic shearing construction in [BZSV24, §6.8]. Let $\mathbb{Z}_{\{q\}}=\mathbb{Z}[q^{-1}]$.
@@ 3.3.4. Integral structure
Let $\rho$ (resp. $\rho_{(X)}$, $\rho_{P(X)}$) be the half sum of positive roots for $\mathbf{G}$ (resp. $\mathbf{L}(\mathbf{X})$, $\mathbf{P}(\mathbf{X})$). Define $\fatslash\mathbb{Z}_{\{q\}}[[A^{*}]]^{{\mathbin{\mkern-6.0mu\fatslash}}}$ to be the functions in $\mathbb{C}[[A^{*}]]$ of the form
@@
where $a_{\check{\lambda}}\in\mathbb{Z}_{\{q\}}$. We are again identifying $\mathbb{C}[[A^{*}]]$ with $\mathbb{C}[[\Lambda]]$ by (1).
The ring $\fatslash\mathbb{Z}_{\{q\}}[[\delta_{(X)}^{\frac{1}{2}}A_{X}^{*}]]^{{\mathbin{\mkern-6.0mu\fatslash}}}$ consists of restrictions of functions in $\fatslash\mathbb{Z}[[A^{*}]]^{{\mathbin{\mkern-6.0mu\fatslash}}}$.
Lemma 3.10.
Let $\mathcal{L}$ be a Hecke operator such that $\fatslash\hat{\mathcal{L}}_{(X)}\in\mathbb{Z}_{\{q\}}[\delta_{(X)}^{\frac{1}{2}}A_{X}^{*}]^{{\mathbin{\mkern-6.0mu\fatslash}}}$, in the notation of (6). Let $\phi=\mathcal{L}\cdot\Phi_{0}$, then $\phi(x)\in\mathbb{Z}[q^{-1}]$ for all $x\in X$.
Proof.
Under the various hypotheses discussed in the previous section, the second term in Theorem 3.6 and Conjecture 3.8 lies in the ring $\fatslash\mathbb{Z}_{\{q\}}[[\delta^{\frac{1}{2}}_{(X)}A_{X}^{*}]]^{\mathbin{\mkern-6.0mu\fatslash}}$. It follows that
$\phi(x_{\check{\lambda}})$ $\in\delta_{P(X)}^{-\frac{1}{2}}(x_{\check{\lambda}})\delta_{(X)}^{\frac{1}{2}}(x_{\check{\lambda}})q^{\langle\check{\lambda},\rho\rangle}\mathbb{Z}_{\{q\}}$
$=q^{\langle\check{\lambda},-\rho_{P(X)}+\rho_{(X)}+\rho\rangle}\mathbb{Z}_{\{q\}}$
Since $\rho=\rho_{P(X)}+\rho_{(X)}$ and $2\rho_{P(X)}$ is integral, this is an integral power of $q$. ∎
3.4. Divisibility properties
For Euler system applications, we are interested in proving that for certain $x$, the values $\phi(x)$ are in fact divisible by q1q-1. By the previous calculations, the values $\phi(x)$ are polynomials in $\mathbb{Z}[q^{\pm 1}]$. Now treat $q$ as a formal variable. To prove divisibility by q1q-1, it remains to specialize at $q=1$ and prove that the value of this polynomial is 0. We will do this by an anti-symmetry argument.
3.4.1. Strongly tempered case
Let $X$ be a strongly tempered spherical variety. In [BZSV24, §4.3, 4.4], the authors attached a symplectic $\check{G}$-representation to $X$
$S_{X}=S_{X}^{\circ}\oplus S_{X}^{\bullet},\quad S_{X}^{\bullet}\simeq(S_{X}^{\circ})^{\ast}$
By [BZSV24, Proposition 9.3.3], the multiset of weights of $S_{X}$ exactly agrees with $\Theta_{X}:=\Theta_{X}^{+}\sqcup(-\Theta_{X}^{+})$ under the assumptions of Theorem 3.6 or Conjecture 3.8. While we are citing general results, in each individual case, the fact that $\Theta$ is exactly the multiset of weights of a symplectic representation can be checked explicitly. Moreover, the assumptions also imply that the weights appearing in $S_{X}$ are all minuscule.
According to the polarization, we have another decomposition $\Theta=\Theta^{\circ}\sqcup\Theta^{\bullet}$. Let $\Theta^{\circ,+}=\Theta^{\circ}\cap\Theta^{+}$. This is the multiset of positive weights occurring in $S_{X}^{\circ}$, where positive means lying in the cone $\mathcal{C}_{X}$ defined in Remark 3.7. Note that we again have $\Theta^{\circ}=\Theta^{\circ,+}\sqcup(-\Theta^{\circ,+})$.
Let
$\hat{\mathcal{L}}=\prod_{\check{\theta}\in\Theta_{X}^{\circ}}(1-q^{-\frac{1}{2}}e^{\check{\theta}})\in\mathbb{C}[\Lambda]$
By construction, this is $W$-invariant, so it is the image of a Hecke operator $\mathcal{L}\in\mathcal{H}(G)$. Since $\Theta_{X}$ consists of minuscule weights, this lies in the integral ring $\fatslash\mathbb{Z}_{\{q\}}[A^{*}]^{\mathbin{\mkern-6.0mu\fatslash}}$. If $S_{X}^{\circ}$ is irreducible, then $\mathcal{L}$ is exactly the Hecke polynomial attached to the representation $S_{X}^{\circ}$ evaluated at $q^{-\frac{1}{2}}$.
@@ 3.4.1. Strongly tempered case
Let $\mathcal{L}$ be the Hecke operator defined above, and let $\phi=\mathcal{L}\cdot\Phi_{0}$. If $x$ lies on a wall of type T, then $\phi(x)\in(q-1)\mathbb{Z}[q^{-1}]$.
@@
By the inverse Satake transform, the value $\phi(x_{\check{\lambda}})|_{q=1}$ is the coefficient of $e^{\check{\lambda}}$ in the power series expansion of
$\hat{\mathcal{L}}|_{q=1}\cdot\frac{\mathtt{ad}}{\mathtt{L}},\quad\text{where }\mathtt{ad}=\prod_{\check{\gamma}\in\check{\Phi}_{X}^{+}}(1-e^{\check{\gamma}}),\quad\mathtt{L}=\prod_{\check{\theta}\in\Theta_{X}^{+}}(1-e^{\check{\theta}}).$
For our choice of $\mathcal{L}$, this expression simplifies to
$\mathtt{ad}\cdot\frac{\prod_{\check{\theta}\in\Theta_{X}^{\circ}}(1-e^{\check{\theta}})}{\prod_{\check{\theta}\in\Theta_{X}^{+}}(1-e^{\check{\theta}})}$ $=\mathtt{ad}\cdot\prod_{\check{\theta}\in\Theta_{X}^{\circ,+}}\frac{1-e^{-\check{\theta}}}{1-e^{\check{\theta}}}$
$=\mathtt{ad}\cdot\prod_{\check{\theta}\in\Theta_{X}^{\circ,+}}(-e^{-\check{\theta}})$
This is a polynomial in $\mathbb{C}[\Lambda_{X}]$, so the subtlety explained in Remark 3.7 no longer applies. Denote the second term above by $\mathtt{L}^{\prime}$.
Suppose $\gamma$ is a spherical root of type T, with corresponding coroot $\check{\gamma}$. Let $w_{\gamma}\in W_{X}$ be its associated simple reflection. We consider the behaviour of the above expression under $w_{\gamma}$. There are two terms:
@@
    We need to understand the difference between $w_{\gamma}\Theta_{X}^{+}$ and $\Theta_{X}^{+}$. In the homogeneous case, this is described in [Sak13, Statement 7.1.5]. Since $\gamma$ has type T, there are two virtual colors $D$ and $D^{\prime}$ belonging to $\gamma$. Their associated valuations $v_{D}$ and $v_{D^{\prime}}$ sum to $\check{\gamma}$, so
    $\frac{w_{\gamma}\mathtt{L}^{\prime}}{\mathtt{L}^{\prime}}=e^{-(v_{D}+v_{D^{\prime}})}=e^{-\check{\gamma}}$
    In general, the above discussion still holds due to our minuscule hypothesis combined with [SW22, Theorem 7.1.9(iii)].
@@
Performing the above computation for a spherical root of type G recovers equation (7), except with a plus sign. Therefore, we gain no information. In particular, all roots are of type G in the group case $\mathbf{H}\backslash\mathbf{H}\times\mathbf{H}$, so we do not see any automatic divisibility in the classical Satake isomorphism.
Remark 3.13.
It is important that we reduced to a polynomial first before considering the action of the Weyl group, since a reflection $w_{\gamma}$ changes the cone $\mathcal{C}_{X}$ that we are expanding along, and there is no a priori relation between the coefficients of the two expansions. Note that this cancellation happened exactly because of our choice of $\mathcal{L}$. This fits nicely with the optimality results of A. Groutides, cf. [Gro24].
3.4.2. Non-strongly tempered case — an example
In the general case, the structure of the denominator cancellation is more subtle due to the extra term in [BZSV24, (4.16)]. Since we only have one application in mind which is not strongly tempered, namely the Friedberg–Jacquet setting, we will write down everything explicitly and observe the cancellation “by hand”. This is also an opportunity to give an example.
In this subsection, we specialize to the following setting:
$\mathbf{H}=\mathrm{GL}_{n}\times\mathrm{GL}_{n}\hookrightarrow\mathbf{G}=\mathrm{GL}_{2n}\times\mathbb{G}_{m},\quad(h_{1},h_{2})\mapsto\Big(\operatorname{diag}(h_{1},h_{2}),\frac{\det h_{1}}{\det h_{2}}\Big)$
The dual group in this case is $\check{G}_{X}=\mathrm{Sp}_{2n}\times\mathbb{G}_{m}$, so it is not strongly tempered. Note that $\mathbf{G}$ contains the torus factor, so in later sections it will be denoted by $\widetilde{\mathbf{G}}$. We avoid the tilde here to simplify notations.
In a suitable coordinate, we have
$\check{\Phi}_{X}^{+}=\{\epsilon_{i}\pm\epsilon_{j}\,|\,1\leq i<j\leq n\}\cup\{2\epsilon_{i}\,|\,1\leq i\leq n\}$
The spherical coroots consist of $\{\epsilon_{i}-\epsilon_{i+1}\,|\,1\leq i<n\}$ and $2\epsilon_{n}$. Only the final one is of type T.
Following the recipe of [Sak13, §7.1], one can compute that
$\Theta_{X}^{+}=\{(\epsilon_{i}\pm\epsilon_{j},2)\,|\,1\leq i<j\leq n\}\cup\{(\epsilon_{i}\pm s,1)\,|\,1\leq i\leq n\},$
where the notation $s$ for the cocharacter on $\mathbb{G}_{m}$ is chosen in reference to its ultimate role as the $s$-variable in the $L$-function. In fact, in this interpretation, the first set corresponds to $L(1,\wedge^{2},\pi)$, and the second term corresponds to the product $L(\frac{1}{2}+s,\pi)L(\frac{1}{2}-s,\pi)$.
In the above coordinates, the Hecke operator giving rise to the standard degree $2n$ $L$-function has the Satake transform
$\hat{\mathcal{L}}=\prod_{i=1}^{n}\Big(1-q^{-\frac{1}{2}}e^{-\epsilon_{i}+s}\Big)\Big(1-q^{-\frac{1}{2}}e^{\epsilon_{i}+s}\Big)$
We now prove exactly the same statement as Proposition 3.11.
Proposition 3.14.
Let $\mathcal{L}$ be the Hecke operator defined above, and let $\phi=\mathcal{L}\cdot\Phi_{0}$. If $x$ lies on a wall of type T, then $\phi(x)\in(q-1)\mathbb{Z}[q^{-1}]$.
Proof.
As before, the value $\phi(x_{\check{\lambda}})|_{q=1}$ is given by the coefficient of $e^{\check{\lambda}}$ in the power series expansion of the expression
$\hat{\mathcal{L}}|_{q=1}\cdot\frac{\prod_{\check{\gamma}\in\check{\Phi}_{X}^{+}}(1-e^{\check{\gamma}})}{\prod_{\check{\theta}\in\Theta_{X}^{+}}(1-e^{\check{\theta}})}.$
By comparing the above descriptions of the multisets $\check{\Phi}_{X}^{+}$ and $\Theta_{X}^{+}$, we see that most terms of the fraction cancels, and it simplifies to
$\prod_{i=1}^{n}(1-e^{-\epsilon_{i}+s})(1-e^{\epsilon_{i}+s})\cdot\prod_{i=1}^{n}\frac{1-e^{2\epsilon_{i}}}{(1-e^{\epsilon_{i}+s})(1-e^{\epsilon_{i}-s})}$ $=\prod_{i=1}^{n}(1-e^{2\epsilon_{i}})\cdot\frac{1-e^{-\epsilon_{i}+s}}{1-e^{\epsilon_{i}-s}}$
$=\prod_{i=1}^{n}(1-e^{2\epsilon_{i}})(-e^{-\epsilon_{i}+s})$
The only root of type T corresponds to the reflection $\epsilon_{n}\mapsto-\epsilon_{n}$. This clearly multiplies the final expression by 1-1. The proposition follows. ∎
3.4.3. Non-split case — an example
When we study the twisted Friedberg–Jacquet setting, we will also encounter a case where $\mathbf{G}$ is not split over $F$. This is not covered by [Sak13], but the necessary inversion formula can be computed using the same method(See the Appendix for details). The result strongly resembles Theorem 3.6, and we will define the combinatorial data appearing in it by analogy.
Let $E/F$ be the unramified quadratic extension with associated quadratic character $\eta$ and non-trivial involution $\bar{\cdot}$. Let $w\in\mathrm{GL}_{n}(E)$ be the anti-diagonal matrix with entities $1$. Define
$\mathrm{U}_{2n}=\{g\in\mathrm{GL}_{2n}(E)|\bar{g}^{T}\begin{pmatrix}0&w\\ w&0\end{pmatrix}g=\begin{pmatrix}0&w\\ w&0\end{pmatrix}\}$.
This is the quasi-split unitary group in $2n$ variables. Consider the following embedding
$\mathbf{H}=\operatorname{Res}_{E/F}\mathrm{GL}_{n}\hookrightarrow\mathbf{G}=\mathrm{U}_{2n}\times\mathbb{G}_{m}\quad h\mapsto(\operatorname{diag}(h,w\bar{h}^{T}w),\mathrm{Norm}_{F}^{E}\det h)$
As in the previous subsection, the group $\mathbf{G}$ here will play the role of $\widetilde{\mathbf{G}}$ in later sections. The group $\mathbf{G}$ is not split, and we recall that in this case, the Satake isomorphism is
[…]
where $W$ is the Weyl group for the root system of type $\mathrm{C}_{n}$.
Set $\mathbf{X}:=\mathbf{G}/\mathbf{H}$. This is a form of the variety $\mathbf{X}$ from the previous subsection. Let
$\Lambda_{X}^{+}=\{(\lambda_{1},\cdots,\lambda_{n})\in\mathbb{Z}^{n}\,|\,\lambda_{1}\geq\cdots\geq\lambda_{n}\}.$
Then we have the generalized Cartan decomposition
$X=\bigsqcup_{\check{\lambda}\in\Lambda_{X}^{+}}x_{\check{\lambda}}K,$
which is analogous to Proposition 3.2. We define the spherical root system to have positive coroots
$\check{\Phi}_{X}^{+}=\{\epsilon_{i}\pm\epsilon_{j}\,|\,1\leq i<j\leq n\}\cup\{2\epsilon_{i}\,|\,1\leq i\leq n\}.$
It is a root system of $\mathrm{C}_{n}$. Among the (simple) spherical roots, we call $2\epsilon_{n}$ the only root of type T. This is still in agreement with the previous subsection when base changed to $E$.
The analogue of $\Theta_{X}^{+}$ now contains the additional data of signs $\sigma_{\check{\theta}}\in\{\pm 1\}$. This was also present in [Sak13] to account for non-split $\mathbf{H}$. Define
$\Theta_{X}^{+}=\{(\epsilon_{i}\pm\epsilon_{j},-,2)\,|\,1\leq i<j\leq n\}\cup\{(\epsilon_{i}\pm 2s,-,2)\,|\,1\leq i\leq n\}.$
Then the analogue of Theorems 3.5 and 3.6 in our case is the following result, whose proof will be given in the appendix (Theorem A.10).
Theorem 3.15.
  1. (1)
[…]
    $C_{c}^{\infty}(X,\mathbb{C})^{K}\cong\mathbb{C}[q^{\pm z_{1}},\cdots,q^{\pm z_{n}},q^{\pm s}]^{W}$
[…]
  2. (2)
    The isomorphism sends the basic function $\mathbf{1}[\mathbf{X}(\mathcal{O})]$ to the constant function 1.
  3. (3)
    There is an inversion formula
    $\delta_{P(X)}^{\frac{1}{2}}\cdot(\mathcal{L}\cdot\Phi_{0})=\hat{\mathcal{L}}\cdot\frac{\prod_{\check{\gamma}\in\check{\Phi}_{X}^{+}}(1-e^{\check{\gamma}})}{\prod_{(\check{\theta},r_{\check{\theta}},d_{\check{\theta}})\in\Theta_{X}^{+}}(1-r_{\check{\theta}}q^{-\frac{1}{2}d_{\check{\theta}}}e^{\check{\theta}})}\bigg|_{\Lambda_{X}^{+}}$
    with
    $\delta_{P(X)}^{1/2}(x_{\lambda})=(-1)^{\sum_{i}\lambda_{i}(n-i+1)}q^{\sum_{i}\lambda_{i}(n-i+1/2)}$
The Hecke operator giving rise to the degree $4n$ $L$-function has the Satake transform
$\hat{\mathcal{L}}_{\mathrm{BC}}$ $=\prod_{i=1}^{n}\Big(1-qe^{-2\epsilon_{i}+2s}\Big)\Big(1-qe^{2\epsilon_{i}+2s}\Big)$
$=\prod_{i=1}^{n}\Big(1-q^{1/2}e^{-\epsilon_{i}+s}\Big)\Big(1-q^{-\frac{1}{2}}e^{\epsilon_{i}+s}\Big)\prod_{i=1}^{n}\Big(1+q^{1/2}e^{-\epsilon_{i}+s}\Big)\Big(1+q^{-\frac{1}{2}}e^{\epsilon_{i}+s}\Big)$
According to this decomposition, write $\hat{\mathcal{L}}_{\mathrm{BC}}=\hat{\mathcal{L}}_{\emptyset}\cdot\hat{\mathcal{L}}_{\eta}$ As in Proposition 3.14, one finds that
$\hat{\mathcal{L}}_{\eta}|_{q=1}\cdot\frac{\prod_{\check{\gamma}\in\check{\Phi}_{X}^{+}}(1-e^{\check{\gamma}})}{\prod_{\check{\theta}\in\Theta_{X}^{+}}(1-e^{\check{\theta}})}.$
simplifies to
$\prod_{i=1}^{n}(1+e^{-\epsilon_{i}+s})(1+e^{\epsilon_{i}+s})\cdot\prod_{i=1}^{n}\frac{1-e^{2\epsilon_{i}}}{(1+e^{\epsilon_{i}+s})(1+e^{\epsilon_{i}-s})}=\prod_{i=1}^{n}(1-e^{2\epsilon_{i}})(e^{-\epsilon_{i}+s})$
From this, exactly the same anti-symmetry argument from before gives the following result.
Proposition 3.16.
Let $\mathcal{L}$ be the Hecke operator whose Satake transform is $\hat{\mathcal{L}}_{\mathrm{BC}}$, and let $\phi=\mathcal{L}\cdot\Phi_{0}$. If $x$ lies on a wall of type T, then $\phi(x)\in(q-1)\mathbb{Z}[q^{-1}]$.